OpenStudy (anonymous):

use the method of substitution and solve these two equations 2x-y=3 7x-3y=11 I got an answer but it just doesn't look right

OpenStudy (ajprincess):

Can u plz post ur workings?

OpenStudy (anonymous):

7x-3(3-2x)=11 7x-9-6x=11 7x+9-9-6x=11+9 x=21 y=3-2x

OpenStudy (anonymous):

y=3-2(21) y=-42

OpenStudy (ajprincess):

2x-y=3 2x-3=y Nw substitute this and u will get t right.

OpenStudy (anonymous):

so the answer would be (21,39)

OpenStudy (ajprincess):